The length (In words) of seven aricles rom a particular newspeper are isted below. Find the mean, median, and mode of the dats, if possible. any of hese measures cannot be found or a measure does not represent the center of the dats expiain why 816 1276 1227 1369 1277 12 154 Find the mean word count Select the correet chaice below and, if necessary, i in he answer box to complete your choice. O A The mean word count is (Round toore dermal pace อเs needed ) B. There is no mean甽count. Does the mean represent the center of the data? O A The mean represents the center. B. The mean does not represent tho centior because it is the smalest data value. O C. The mean does not represent the center because it is the largest data value. OD, The mean does not represent the center because is rot a data volue. O E. There is no mean word count Find the madian wond count. Select the correct cholice below and, f necessary,l in te answer bax to complete your choice O A The median word count is Round to one deaimal place as needed) O a. There is no medan wond counmt Does the medan represent Pe ㎝rter of the data? O A. The median represents the center B The median does not represent the center because it is the smallest data value. O C. The median does not represent the center because it is not a data value. O D. The median does not represent the center because it is the largest cata value O t. There is no medan wordcount Find the mode of the word couts Select tho correct choice wow and, d recessary,個ite arawer box to complete your doce Click to select your answer
Find the mode of the word counts. Select the comect choice below and, if necessary, Sitl in the answer box to complete your choice O A The mode/s) of the word counts is (are) Round to one decimal place as needed. Use a comma to separate answers as needed) O B. There is no mode. Does (Do) the mode(s) represent the center of the data? 。A. The mode(s) represent(s) the center O B. The mode(s) does (do) not represent the center because it (one) is the largest data value 。c. The mode(s) does (de) nol represent he center because (ney) is (are) not a data vale ○D. The modees) does (do)noe nesresent the center becane it(ore-Pe sn ales' data value OE· There is no mode Cick to select your answer
Answer #1

(a)

Mean = 8247/7 = 1178.1

Correct option:

A.   The mean word count is 1178.1

(b)

Correct option:

A. The mean represents the center.

(c)

Arranging data in ascending order, we get:

816, 1128, 1154, 1227, 1276, 1277, 1369

n = 7

Median is (7 = 1)/2th item = 4th item = 1227

So,

median = 1227

So,

Correct option :

A.. The median word count is : 1227

(d)

Correct option:

A. The median represents the center.

(e)

Since each value occurs only one, there is no mode for this data set.

Correct option:

B.. There is no mode

(f)

Correct option:

E. There is no mode